OpenStudy (usernamenottaken):

Will Medal! Pre Calc help If csc Θ=0.8 and 270°<Θ<360°, find the exact value of sin 2Θ.

OpenStudy (anonymous):

Hi @UsernameNotTaken ready to solve it?

OpenStudy (astrophysics):

ok

OpenStudy (anonymous):

@Astrophysics s/he should know that this problem is incorrect. shouldn't s/he?

OpenStudy (usernamenottaken):

I was afk doing other problems, sorry for not being here in time.

OpenStudy (usernamenottaken):

Yes i did make a mistake... The correct version is cos Θ=0.8 and 270°<Θ<360°, find the exact value of sin 2Θ.

OpenStudy (astrophysics):

Ok better, so \[\sin 2 \theta = 2 \sin \theta \cos \theta \]

OpenStudy (anonymous):

:)

OpenStudy (usernamenottaken):

xD yeah sorry, he he he...

OpenStudy (astrophysics):

Do you know how to find sin theta?

OpenStudy (usernamenottaken):

my answer is -0.96 but i think its wrong..

OpenStudy (usernamenottaken):

@astrophysics yes i do

OpenStudy (astrophysics):

I think your answer maybe ok, lets keep going and see if it's right or not

OpenStudy (astrophysics):

So what did you get for sin theta?

OpenStudy (anonymous):

Why do you think so? Explain how you solved it and @Astrophysics and me can check it. ur answer's correct.

OpenStudy (astrophysics):

Thanks @Yavar

OpenStudy (usernamenottaken):

270° < θ < 360°, θ is in Quadrant III. so sinθ < 0. By drawing a 3-4-5 triangle in Quadrant III such that cosθ = 3/5, we see that: adj = 3, opp = 4, and hyp = 5 ==> sinθ = -opp/hyp = -4/5. Therefore, by the double-angle formula for sine: sin(2θ) = 2sinθcosθ = 2(-4/5)(3/5) = -24/25. or -0.96

OpenStudy (usernamenottaken):

That was basically it in a nutshell....

OpenStudy (astrophysics):

Hey, that looks good!

OpenStudy (anonymous):

That was nice.....

OpenStudy (triciaal):